How long will it take to have R30 000 for a trip to Europe if I invest R20 000 now at 15% per year,compounded quarterly?
Answer by Alan3354(69443)

How long will it take to have R30 000 for a trip to Europe if I invest R20 000 now at 15% per year,compounded quarterly?
=================
15%/yr = 3.75%/quarter ---> 0.0375
--------
30000 = 20000*(1 = 0.0375)^t (t in quarters)
3 = 2*1.0375^t
1.0375^t = 3/2
t*log(1.0375) = log(1.5)
t = log(1.5)/log(1.0375)
t =~ 11 quarters